# Explaining Text Classifiers with Counterfactual Representations

### v2

@@ 5.2 Treatment effect on synthetic data
Both $\widehat{\mathrm{ATE}}_{\widehat{Y}}$ and $\widehat{\mathrm{TE}}_{\widehat{Y}}$ should be thought of as estimators for corresponding quantities ${\mathrm{ATE}}_{\widehat{Y}}$ and ${\mathrm{TE}}_{\widehat{Y}}$ defined just as in (12) except that the representations $X(s_{Z\leftarrow z})$ of the true CFs are substituted for the CFR $X(s)_{Z\leftarrow z}$.
@@ 5.3 Treatment effect on realistic data
$\widehat{\mathrm{ATE}}^{\mathrm{score}}_{\widehat{Y}}[\mathcal{S}^{(z_{1},z_{2})}]:=\frac{1}{|\mathcal{S}^{(z_{1},z_{2})}|}\sum_{s\in\mathcal{S}^{(z_{1},z_{2})}}\Big{(}\widehat{Y}(X(s)_{Z\leftarrow z})-\widehat{Y}(X(s))\Big{)}$ (13)
@@
To assess how well CFRs account for individual causal effects and also having various approaches to explainability in mind, we adapt the error measure introduced in (Abraham et al. (2022), definition 3) to our CFRs:
$\mathrm{Error}_{\widehat{Y}}[\mathcal{S}]:=\frac{1}{|\mathcal{S}|}\sum_{(s,z)\in\mathcal{S}}\mathrm{Dist}\Big{(}\>p_{\widehat{Y}}\big{(}X(s_{Z\leftarrow z})\big{)}\!-\!p_{\widehat{Y}}\big{(}X(s)\big{)}\>,\>p_{\widehat{Y}}\big{(}X(s)_{Z\leftarrow z}\big{)}\!-\!p_{\widehat{Y}}\big{(}X(s)\big{)}\>\Big{)}$ (14)
where Dist is a distance between the observed individual effects and the individual effects estimated using CFRs. Following Abraham et al. (2022), we consider three distance measures: the cosine distance which is influenced only by the directions of the effects, the normdiff which is the absolute difference between the Euclidian norms of each effect and is influenced only by the magnitude of the effects and at last the $L^{2}$ distance which is the norm of the difference of effects and is influenced by both the magnitude and direction of the effects [^6] Comparison of individual effects in section 5.2 is based on the distance in total variation rather than the Euclidian norm of the effects, which in both cases amounts to considering magnitude only..
@@ Results
Table 4: $\mathrm{Error}_{\widehat{Y}}$ (and standard deviations) for a 5-way sentiment linear classifier on top of a frozen finetuned bert-base-uncased previsously finetuned for this task. Rows are distances. Columns are explanatory methods. Lower is better. Best result per metric are highlighted in bold. Results are averaged over 10 random initializations. The random explainer takes the difference between two random probability vectors as the predicted effect.
@@ 5.4 Explaining predictions on real-world data
In this subsection we investigate the usefulness of our CFRs as a practical tool for providing a detailed bias analysis in a real world example. We select the BiasInBios dataset (De-Arteaga et al., 2019) because it is notoriously biased and approximate genuine counterfactuals can be generated from the observations. The authors create these CFs by simply swapping the gender $z$ for its opposite $\bar{z}$ in each biography [^7] These CFs are flawed because other factors correlated with gender are not modified by swapping gender indicators.. These CFs thus provides us with a ground truth against which we can compare our CFRs.
The bias analysis in De-Arteaga et al. (2019) revolves around a somewhat involved miss-classification rate that we now embark to adapt for our purpose. Let’s fix a gender $z$ and two occupations $y_{\mathrm{f}}\neq y_{\mathrm{t}}$. First consider the subset of sentences $s$ with a gender $z$ which are misclassified as $y_{\mathrm{f}}$ when the classifier $\widehat{Y}$ uses the original representation $X(s)$ while it makes a correct prediction $y_{\mathrm{t}}$ when using the CFR $X(s)_{Z\leftarrow\bar{z}}$ for the swapped gender $\bar{z}$. We next consider the larger subset where we relax the misclassification constraint. We then define a missclassification rate $\widehat{\Pi}_{z,(y_{\mathrm{f}},y_{\mathrm{t}})}$ as a ratio between the cardinalities of the former to the latter
@@
This missclassification rate $\widehat{\Pi}_{z,(y_{\mathrm{f}},y_{\mathrm{t}})}$ can be thought of as an estimator for a quantity $\Pi_{z,(y_{\mathrm{f}},y_{\mathrm{t}})}$ defined as in (15) except that the representations $X(s_{Z\leftarrow\bar{z}})$ of the genuine CFs are used instead of the CFRs. At last we define $\widehat{\Pi}^{\text{max}}_{z}$ as the maximum of $\widehat{\Pi}_{z,(y_{\mathrm{f}},y_{\mathrm{t}})}$ over all possible pairs $(y_{\mathrm{f}},y_{\mathrm{t}})$.
@@ 5.5 CFRs beyond explainability
Part of the usefulness of our CFRs stems from the possibility to compute them even in circumstances where no explicit text CF would make sense. However, as a simple consistency check, it is tempting to ask how CFR work on single word representations such as GloVe embeddings, which are notoriously gender-biased (Bolukbasi et al. (2016); Ravfogel et al. (2022)). More precisely, for a word $s$ with a given $Z(s)=z$ we can ask which word $s^{\prime}$ has the closest embedding $X(s^{\prime})$ to the CFR $X(s)_{Z\leftarrow z^{\prime}}$, thus providing an explicit approximate textual counterfactual. We performed many such checks in Supplementary material C when $Z$ corresponds to a gender bias. For example the word $s=$bridesmaids" becomes $s^{\prime}=$groomsmen" through such an indirect gender switch. Most examples are indeed convincing explicit gender counterfactuals.
@@ B.1 Treatment effect on EEEC+
Figure 3: Evolution for aggressive training scenarios of $\mathrm{ATE}_{\widehat{Y}}[\mathcal{S}_{n}]$ and $\widehat{\mathrm{ATE}}_{\widehat{Y}}[\mathcal{S}_{n}]$ (top) of the correlation coefficient $\rho$ (middle) and the linear regression coefficient $\alpha$ (bottom) between $\mathrm{TE}_{\widehat{Y}}$ and $\widehat{\mathrm{TE}}_{\widehat{Y}}$ in $\mathcal{S}_{n}$ vs. the fraction $|\mathcal{S}_{n}|/|\mathcal{S}|$ (in %) of included observations. The dotted vertical lines corresponds to a maximal fraction of observations above which the correlation $\rho$ falls below $0.75$ and $0.5$ respectively.
@@ B.2 Treatment effect on CEBaB
In Table 11 we report the treatment effects $\mathrm{ATE}^{\mathrm{score}}$ and its estimation $\widehat{\mathrm{ATE}}^{\mathrm{score}}$ using CFRs (with binary and ternary settings) or approximate CFs as substitutes for genuine CFs.
@@ Results
This evaluations provide further evidence for the practical usefulness of our CFRs for data augmentation purposes and, indirectly, inspire confidence in their quality.

### v3

@@ 5.2 Treatment effect on synthetic data
Both $\widehat{\mathrm{ATE}}_{\widehat{Y}}$ and $\widehat{\mathrm{TE}}_{\widehat{Y}}$ should be thought of as estimators for corresponding quantities ${\mathrm{ATE}}_{\widehat{Y}}$ and ${\mathrm{TE}}_{\widehat{Y}}$ defined just as in (12) except that the representations $X(s_{Z\leftarrow z})$ of the true CFs are substituted for the CFRs $X(s)_{Z\leftarrow z}$.
@@ 5.3 Treatment effect on realistic data
$\widehat{\mathrm{ATE}}^{\mathrm{score}}_{\widehat{Y}}[\mathcal{S}^{(z_{1},z_{2})}]:=\frac{1}{|\mathcal{S}^{(z_{1},z_{2})}|}\sum_{s\in\mathcal{S}^{(z_{1},z_{2})}}\Big{(}\widehat{Y}(X(s)_{Z\leftarrow z})-\widehat{Y}(X(s))\Big{)}.$ (13)
@@
To assess how well CFRs account for individual causal effects and also having various approaches to explainability in mind, we adapt the error measure introduced in Abraham et al. (2022) (definition 3) to our CFRs:
$\mathrm{Error}_{\widehat{Y}}[\mathcal{S}]=\frac{1}{|\mathcal{S}|}\sum_{(s,z)\in\mathcal{S}}\!\!\!\mathrm{Dist}\Big{(}\>p_{\widehat{Y}}\big{(}X(s_{Z\leftarrow z})\big{)}\!-\!p_{\widehat{Y}}\big{(}X(s)\big{)}\>,\>p_{\widehat{Y}}\big{(}X(s)_{Z\leftarrow z}\big{)}\!-\!p_{\widehat{Y}}\big{(}X(s)\big{)}\>\Big{)}.$ (14)
where Dist is a distance between the observed individual effects and the individual effects estimated using CFRs. Following Abraham et al. (2022), we consider three distance measures: the cosine distance which is influenced only by the directions of the effects, the normdiff which is the absolute difference between the Euclidian norms of each effect and is influenced only by the magnitude of the effects and at last the $L^{2}$ distance which is the norm of the difference of effects and is influenced by both the magnitude and direction of the effects. [^6] Comparison of individual effects in section 5.2 is based on the distance in total variation rather than the Euclidian norm of the effects, which in both cases amounts to considering magnitude only.
@@ Results
Table 4: $\mathrm{Error}_{\widehat{Y}}$ (and standard deviations) for a 5-way sentiment linear classifier on top of a frozen bert-base-uncased previously finetuned for this task. Rows are distances. Columns are explanatory methods. Lower is better. Best results per metric are highlighted in bold. Results are averaged over 10 random initializations. The random explainer takes the difference between two random probability vectors as the predicted effect.
@@ 5.4 Explaining predictions on real-world data
In this subsection we investigate the usefulness of our CFRs as a practical tool for providing a detailed bias analysis in a real world example. We select the BiasInBios dataset (De-Arteaga et al., 2019) because it is notoriously biased and approximate genuine counterfactuals can be generated from the observations. De-Arteaga et al. (2019) create these CFs by simply swapping the gender $z$ for its opposite $\bar{z}$ in each biography [^7] These CFs are flawed because other factors correlated with gender are not modified by swapping gender indicators.. These CFs thus provides us with a ground truth against which we can compare our CFRs.
The bias analysis in De-Arteaga et al. (2019) revolves around a somewhat involved miss-classification rate that we now embark to adapt for our purpose. Let’s fix a gender $z$ and two occupations $y_{\mathrm{f}}\neq y_{\mathrm{t}}$. First consider the subset of sentences $s$ with a gender $z$ which are misclassified as $y_{\mathrm{f}}$ when the classifier $\widehat{Y}$ uses the original representation $X(s)$ while it makes a correct prediction $y_{\mathrm{t}}$ when using the CFR $X(s)_{Z\leftarrow\bar{z}}$ for the swapped gender $\bar{z}$. We next consider the larger subset where we relax the misclassification constraint. We then define a misclassification rate $\widehat{\Pi}_{z,(y_{\mathrm{f}},y_{\mathrm{t}})}$ as a ratio between the cardinalities of the former to the latter
@@
This misclassification rate $\widehat{\Pi}_{z,(y_{\mathrm{f}},y_{\mathrm{t}})}$ can be thought of as an estimator for a quantity $\Pi_{z,(y_{\mathrm{f}},y_{\mathrm{t}})}$ defined as in (15) except that the representations $X(s_{Z\leftarrow\bar{z}})$ of the genuine CFs are used instead of the CFRs. At last we define $\widehat{\Pi}^{\text{max}}_{z}$ as the maximum of $\widehat{\Pi}_{z,(y_{\mathrm{f}},y_{\mathrm{t}})}$ over all possible pairs $(y_{\mathrm{f}},y_{\mathrm{t}})$.
@@ 5.5 CFRs beyond explainability
Part of the usefulness of our CFRs stems from the possibility to compute them even in circumstances where no explicit text CF would make sense. However, as a simple consistency check, it is tempting to ask how CFR work on single word representations such as GloVe embeddings, which are notoriously gender-biased Bolukbasi et al. (2016); Ravfogel et al. (2022). More precisely, for a word $s$ with a given $Z(s)=z$ we can ask which word $s^{\prime}$ has the closest embedding $X(s^{\prime})$ to the CFR $X(s)_{Z\leftarrow z^{\prime}}$, thus providing an explicit approximate textual counterfactual. We performed many such checks in Supplementary Material C when $Z$ corresponds to a gender bias. For example the word $s=$bridesmaids" becomes $s^{\prime}=$groomsmen" through such an indirect gender switch. Most examples are indeed convincing explicit gender counterfactuals.
@@ B.1 Treatment effect on EEEC+
Figure 3: Evolution for aggressive training scenarios of $\mathrm{ATE}_{\widehat{Y}}[\mathcal{S}_{n}]$ and $\widehat{\mathrm{ATE}}_{\widehat{Y}}[\mathcal{S}_{n}]$ (top) of the correlation coefficient $\rho$ (middle) and the linear regression coefficient $\alpha$ (bottom) between $\mathrm{TE}_{\widehat{Y}}$ and $\widehat{\mathrm{TE}}_{\widehat{Y}}$ in $\mathcal{S}_{n}$ vs. the fraction $|\mathcal{S}_{n}|/|\mathcal{S}|$ (in %) of included observations. The dotted vertical lines corresponds to a maximal fraction of observations above which the correlation coefficient $\rho$ falls below $0.75$ and $0.5$ respectively.
@@ B.2 Treatment effect on CEBaB
In Table 11 we report the treatment effect $\mathrm{ATE}^{\mathrm{score}}$ and its estimation $\widehat{\mathrm{ATE}}^{\mathrm{score}}$ using CFRs (with binary and ternary settings) or approximate CFs as substitutes for genuine CFs.
@@ Results
This evaluation provides further evidence for the practical usefulness of our CFRs for data augmentation purposes and, indirectly, inspire confidence in their quality.